1. George runs 10 miles per hour and Bill runs 6 miles per hour. They start on a jogging course at the same time and George finishes 1 hour ahead of Bill. How long is the course?
-----------------------------------
2. Karen takes 5 hours to count her store inventory. When Anna helps, they can count the inventory in 3 hours. How long will it take Anna to count inventory if she works alone?

Answer by lwsshak3(11628)

1. George runs 10 miles per hour and Bill runs 6 miles per hour. They start on a jogging course at the same time and George finishes 1 hour ahead of Bill. How long is the course?
***
let x=length of course
Travel time=distance/speed
x/6-x/10=1
lcd:30
5x-3x=30
2x=30
x=15
length of course=15 miles
-----------------------------------
2. Karen takes 5 hours to count her store inventory. When Anna helps, they can count the inventory in 3 hours. How long will it take Anna to count inventory if she works alone?
***
1/5=karen's work rate
1/x=anna's work rate
1/3=work rate working together
sum of indv work rates=work rate working together
1/5+1/x=1/3
1/x=1/3-1/5=5/15-3/15=2/15
1/x=2/15
x=15/2
x=7.5
How long will it take Anna to count inventory if she works alone? 7.5 hrs